### v1

# Trivialization of the gravitational Green-Schwarz mechanism in the non-relativistic limit of string theory

Abstract
We show that the gravitational Green–Schwarz anomaly cancellation mechanism becomes trivial in the non-relativistic (NR) limit of ten-dimensional heterotic supergravity with four-derivative corrections. We perform a NR expansion of the Kalb–Ramond field and identify the finite Green–Schwarz transformation in this limit, which can be interpreted as a non-covariant $SO(8)$ transformation. We then construct an explicit field redefinition such that the redefined two-form is invariant under this symmetry. This result is compared with the previously reported trivialization of the gauge Green–Schwarz mechanism under the same limit. Both field redefinitions can be implemented simultaneously, and the associated Chern–Simons terms are exact, arising directly from the redefinition structure, and leading to a trivial Bianchi identity. These results imply that anomaly cancellation becomes automatic in the NR regime, extending the consistent formulations of NR heterotic string theories and giving rise to a richer class of topological heterotic backgrounds with simpler thermodynamical properties such as the Wald entropy and the first law.
I Introduction
Non-relativistic (NR) string theories NR1 -NR2 emerged as a rich and consistent extension of string theory, motivated by decoupling limits, dualities, and non-Lorentzian geometries (see Review for a modern review). These theories admit worldsheet formulations with modified symmetries and target spaces governed by generalized Newton–Cartan geometry. In particular, the NR limits of supergravity provide a natural low-energy arena in which the interplay between background geometry, field content, and anomaly cancellation can be studied beyond the relativistic regime. Anomaly cancellation is a cornerstone of string theory consistency. In heterotic string theory, gauge and gravitational anomalies are canceled via the Green-Schwarz mechanism GS , which introduces a nontrivial transformation of the Kalb-Ramond $B$-field under both gauge and local Lorentz transformations. These transformations are essential for consistency and cannot be removed by field redefinitions in a relativistic setting. The structure of the anomaly polynomial imposes rigid constraints, most notably restricting the gauge group to either $\mathrm{SO}(32)$ or $E_{8}\times E_{8}$, due to the requirement that the total anomaly factorizes as $I_{12},\sim(\text{tr}F^{2}-\text{tr}R^{2})^{2}$, a distinctive feature of the heterotic string Gross -Gross2 .
In this work, we demonstrate that the Green-Schwarz mechanism becomes trivial in the NR limit of heterotic supergravity. More precisely, we show that up to four-derivative order (i.e., to first order in $\alpha^{\prime}$), the nontrivial gravitational Green-Schwarz transformation of the $B$-field can be entirely removed by field redefinitions. As a consequence, the redefined two-form $\tilde{b}_{\mu\nu}$ becomes manifestly invariant under $SO(8)$ transformations. Moreover, the modified three-form field strength $\bar{h}_{\mu\nu\rho}$ becomes exact, and its Bianchi identity is trivially satisfied. These results, together with the previously reported gauge Green-Schwarz trivialization Lescano2025 , have profound implications: the anomaly cancellation in heterotic string theory no longer imposes constraints on the gauge group or background topology. The need for anomaly inflow is eliminated, and the associated cohomological conditions are no longer required. This leads to a dramatic relaxation of consistency constraints and allows for compactifications and gauge sectors that would be anomalous in the relativistic theory. Similarly, the Wald entropy and the first law of thermodynamics are simplified under this limit, thanks to the simplification of the Bianchi identity for the three-form curvature. These findings open the door to a broader landscape of consistent NR heterotic theories, including those with unconventional gauge groups, torsional geometries, or non-Kähler compactifications upon considering the NR limit. Furthermore, it provides a new perspective on the role of anomalies in string theory, suggesting that non-Lorentzian limits can fundamentally alter their geometric and topological origin. The four-derivative corrections to the universal action principle of the different formulations of heterotic string theory were historically computed considering three- and four-point scattering amplitudes for the massless states GrossSloan -CN . This method is based on the study of the different types of string interactions through the S-matrix to construct an effective Lagrangian, originally computed by Metsaev and TseytlinMetsaevTseytlin ,
$S_{MT}=\int d^{D}x\sqrt{-g}e^{-2\phi}(L^{(0)}+L^{(1)}_{MT})\,,$ (1)
where $L^{(0)}=R+4\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{12}H_{\mu\nu\lambda}H^{\mu\nu\lambda}$ and
$L^{(1)}_{\rm MT}$ $=$ $-\frac{a+b}{8}\Big{[}R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-\frac{1}{2}H^{\mu\nu\rho}H_{\mu\sigma\lambda}R_{\nu\rho}{}^{\sigma\lambda}+\frac{1}{24}H^{4}$ (2)
$-\frac{1}{8}H^{2}_{\mu\nu}H^{2\mu\nu}\Big{]}+\frac{a-b}{4}H^{\mu\nu\rho}C_{\mu\nu\rho}\,$
corresponds to the first-order $\alpha^{\prime}$-correction with $a$ and $b$ units of $\alpha^{\prime}$. Our compact notation is
$C_{\mu\nu\rho}$ $=$ $w_{[\mu}{}^{AB}\partial_{\nu}w_{\rho]AB}+\frac{2}{3}w_{[\mu}{}^{AB}w_{\nu B}{}^{C}w_{\rho]CA}\,,$ (3)
$H^{2}_{\mu\nu}$ $=$ $H_{\mu}{}^{\rho\sigma}H_{\nu\rho\sigma}\,,\quad H^{2}=H_{\mu\nu\rho}H^{\mu\nu\rho}\,,$ (4)
with $A,B=1,\dots,10$ a flat index and $w_{\mu AB}$ the spin connection given by
$w_{\mu BC}$ $=$ $E_{\mu}{}^{A}\left(-E^{\rho}{}_{[A}{}E^{\nu}{}_{B]}{}\partial_{\rho}E_{\nu C}+E^{\rho}{}_{[A}{}E^{\nu}{}_{C]}{}\partial_{\rho}E_{\nu B}\right.$ (5)
$\left.+E^{\rho}{}_{[B}E^{\nu}{}_{C]}\partial_{\rho}E_{\nu A}\right)\,.$
The bi-parametric form of the Lagrangian (2) was introduced in Tduality . In order to find heterotic $\alpha^{\prime}$-corrections we impose $(a,b)=(-\alpha^{\prime},0)$. For the moment, we are turning off the non-Abelian gauge field $A_{\mu}{}^{i}$ of the heterotic supergravity, which we will include later.
The full action (1) requires a modification of the Lorentz transformation of the B-field at order $\alpha^{\prime}$,
$\delta_{\Lambda}B_{\mu\nu}=-\frac{\alpha^{\prime}}{2}\partial_{[\mu}\Lambda^{AB}w_{\nu]AB}\,.$ (6)
This expression is the so-called Green-Schwarz transformation GS and it is the cornerstone of string theory phenomenology. We can absorb the second line of (2) by deforming the curvature of the B-field
$\bar{H}_{\mu\nu\rho}=3\Big{(}\partial_{[\mu}B_{\nu\rho]}+\frac{\alpha^{\prime}}{2}C_{\mu\nu\rho}\Big{)}\,.$ (7)
The new terms in the curvature of the B-field guarantees that $\bar{H}_{\mu\nu\rho}$ transforms trivially under Lorentz transformations and then $L_{\rm MT}$ is Lorentz invariant when four-derivative corrections are taken into account.
II Main Result: Trivialization of the gravitational Green-Schwarz Mechanism
II.1 $\alpha^{\prime}$-rescaling for the non-relativistic limit
Higher-derivative corrections in heterotic string theory arise from quantum consistency of the worldsheet sigma model, with $\alpha^{\prime}$ serving as the loop-counting parameter in target-space curvature corrections. In the NR limit, the expansion of the worldsheet action around a NR background geometry implies a systematic suppression of gravitational curvature effects. From the sigma model perspective, the worldsheet metric couples to a target-space metric with the expansion
$g_{\mu\nu}=c^{2}\tau_{\mu}^{a}\tau_{\nu}^{b}\eta_{ab}+h_{\mu\nu},$ (8)
where $a=0,1$ is a longitudinal flat index and therefore the dominant $c^{2}$ scaling is carried by the longitudinal Newton-Cartan clock form $\tau_{\mu}^{a}$. At leading order, the divergences of the metric are cancelled by expanding the B-field as
$B_{\mu\nu}=-c^{2}{\tau_{\mu}}^{a}{\tau_{\nu}}^{b}\epsilon_{ab}+b_{\mu\nu}\,.$ (9)
As a result, the pullback of the Riemann tensor to the worldsheet includes components that scale as $R\sim c^{2}$, and therefore a curvature-squared term in the effective action diverges in the $c\rightarrow\infty$ limit (this can also be checked by a straightforward computation of the Riemann squared term). To maintain a finite worldsheet theory that yields a consistent NR target-space action, we rescale the relativistic $\alpha^{\prime}$ parameter in the following way,
$\alpha^{\prime}\quad\longrightarrow\quad\frac{\alpha^{\prime}_{NR}}{c^{2}},$ (10)
which precisely compensates the leading divergence in the two-loop beta function for the background fields. This prescription ensures that curvature-squared terms, such as those arising in the gravitational Green-Schwarz counterterm or in the $H^{4}$-terms, survive the NR limit up to covariant field redefinitions (see HSZ for a discussion on the role of the covariant field redefinitions before taking the NR limit when four-derivative terms are considered).
II.2 Field redefinitions and Green-Schwarz trivialization
We start by considering the NR expansion of ten-dimensional heterotic supergravity, following the formalism developed in NSNS , where the components of the vielbein are
$E_{\mu}{}^{A}=(c\,e_{\mu}{}^{a},e_{\mu}{}^{a^{\prime}})$ (11)
and the inverse components are
$E^{\mu}{}_{A}=(\frac{1}{c}\,e^{\mu}{}_{a},e^{\mu}{}_{a^{\prime}})\,.$ (12)
The gravitational Green-Schwarz mechanism for the b-field is given by
$\delta_{\Lambda}b_{\mu\nu}=-\frac{\alpha^{\prime}_{NR}}{2c^{2}}\partial_{[\mu}\Lambda^{AB}w_{\nu]AB}$ (13)
where the Lorentz parameter splits according to a SO(1,1) parameter $\Lambda_{ab}=\lambda_{M}\epsilon_{ab}$, a $SO(8)$ parameter $\Lambda_{a^{\prime}b^{\prime}}=\lambda_{a^{\prime}b^{\prime}}$ and a boost parameter $\Lambda_{ab^{\prime}}=-\Lambda_{b^{\prime}a}=\frac{1}{c}\lambda_{ab^{\prime}}$.
The only contribution upon considering the NR limit $c\rightarrow\infty$ is given by a transverse Green-Schwarz mechanism of the form,
$\delta_{\lambda}b_{\mu\nu}=-\frac{\alpha^{\prime}_{NR}}{4}\partial_{\mu}\lambda^{a^{\prime}b^{\prime}}\tau_{\nu]}{}^{d}\eta_{de}\tau_{\rho\sigma}{}^{e}e^{\rho}{}_{a^{\prime}}e^{\sigma}{}_{b^{\prime}}\,,$ (14)
where $\tau_{\rho\sigma}{}^{e}=2\partial_{[\rho}\tau_{\sigma]}{}^{e}$ is the intrinsic torsion, and we have simplified the last expression by using the Newton-Cartan relations,
$\tau_{\mu}{}^{a}e^{\mu}{}_{a^{\prime}}$ $=\tau^{\mu}{}_{a}e_{\mu}{}^{a^{\prime}}=0\,,\qquad e_{\mu}{}^{a^{\prime}}e^{\mu}{}_{b^{\prime}}=\delta^{a^{\prime}}_{b^{\prime}}\,,$ (15)
$\tau_{\mu}{}^{a}\tau^{\mu b}$ $=\eta^{ab}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (16)
The transversal Green-Schwarz mechanism (13) can be trivialized by the simple observation that $S_{\nu a^{\prime}b^{\prime}}=\tau_{\rho\sigma}{}^{e}\tau_{\nu}{}^{d}\eta_{de}e^{\rho}{}_{a^{\prime}}e^{\sigma}{}_{b^{\prime}}$ transforms covariantly under $SO(8)$, i.e.,
$\delta S_{\nu a^{\prime}b^{\prime}}=S_{\nu c^{\prime}b^{\prime}}\lambda^{c^{\prime}}{}_{a^{\prime}}+S_{\nu a^{\prime}c^{\prime}}\lambda^{c^{\prime}}{}_{b^{\prime}}\,.$ (17)
The field redefinition that trivializes this mechanism is given by
$\tilde{b}_{\mu\nu}=b_{\mu\nu}-\frac{\alpha^{\prime}_{NR}}{4}w_{[\mu}{}^{a^{\prime}b^{\prime}}\tau_{\nu]}{}^{d}\eta_{de}\tau_{\rho\sigma}{}^{e}e^{\rho}{}_{a^{\prime}}e^{\sigma}{}_{b^{\prime}}\,,$ (18)
which exactly absorbs the non-covariant $SO(8)$ transformation of the b-field. This field redefinition leads to a finite, gauge-invariant three-form curvature $\bar{h}_{\mu\nu\rho}$ which we can define as
$\bar{h}_{\mu\nu\rho}=3\partial_{[\mu}\tilde{b}_{\nu\rho]}=h_{\mu\nu\rho}+C_{\mu\nu\rho}\,,$ (19)
where the Chern-Simons 3-form is given by
$C_{\mu\nu\rho}=-\frac{\alpha^{\prime}_{NR}}{4}\partial_{[\mu}\Big{(}w_{\nu}{}^{a^{\prime}b^{\prime}}\tau_{\rho]}{}^{d}\eta_{de}e^{\epsilon}{}_{a^{\prime}}e^{\sigma}{}_{b^{\prime}}\tau_{\epsilon\sigma}{}^{e}\Big{)}\,.$ (20)
II.3 Comparison with the gauge Green-Schwarz trivialization
For comparing the previous trivialization with the gauge one, let us introduce the gauge field $A_{\mu}{}^{i}$ with expansion $\hat{A}_{\mu}^{i}=c\ {\tau_{\mu}}^{-}\alpha_{-}^{i}+\frac{1}{c}a_{\mu}^{i}$. In this expression, $a_{\mu}^{i}$ is a gauge connection and $\alpha_{-}^{i}$ is a vector. This expansion was analyzed in Lescano2025 , where the trivialization of the gauge Green-Schwarz mechanism was observed  [^1] Another possibility to explore the NR limit of heterotic supergravity is to follow the prescription given in BR . In this case the cancellation of the gauge Green-Schwarz mechanism is not straightforward due to non-covariant gauge transformations in the frame fields. However, since that prescription is T-duality invariant, there should exist field redefinitions connecting both Lescano2025 and BR .. Upon taking the limit $c\rightarrow\infty$ the b-field transforms as,
$\delta_{\lambda}b_{\mu\nu}$ $=$ $-(\partial_{[\mu}\lambda^{i})\alpha_{-i}\tau_{\nu]}^{-}\,,\,.$ (21)
with $\lambda^{i}$ an arbitrary gauge parameter, while the gauge field transform as
$\delta_{\lambda}a_{\mu}^{i}$ $=$ $\partial_{\mu}\lambda^{i}+f^{i}{}_{jk}\lambda^{j}a_{\mu}^{k}\,,$ (22)
$\delta_{\lambda}\alpha_{-}^{i}$ $=$ $f^{i}{}_{jk}\lambda^{j}\alpha_{-}^{k}\,.$ (23)
Therefore, one can redefine the b-field as
$\tilde{b}_{\mu\nu}=b_{\mu\nu}+a_{[\mu|i}\alpha_{-}^{i}\tau_{|\nu]-}$ (24)
which is gauge invariant, i.e. $\delta_{\lambda}\tilde{b}_{\mu\nu}=0$. Since all the terms are $SO(8)$ invariant in the previous expression, both Green-Schwarz mechanisms can be trivialized simultaneously. While the leading order of the gauge Lagrangian was explicitly computed in Lescano2025 , the field redefinition (18) does not guarantee the finiteness of the $L^{1}$ Lagrangian, since covariant field redefinitions may still be necessary for the compensation of potential divergences coming from the Riem2 term.
III Implications
The trivialization of the gravitational and gauge Green-Schwarz mechanism in the NR limit of heterotic supergravity leads to a profound restructuring of the constraints of heterotic string theory. We now outline the main consequences.
III.1 Anomaly cancellation
In relativistic heterotic string theory, one-loop gauge and gravitational anomalies are canceled by a nontrivial transformation of the Kalb-Ramond $B$-field, via its coupling to Chern-Simons terms GS . In the NR limit, the Green-Schwarz transformations are absorbed through field redefinitions as showed in the previous section, and therefore the redefined $\tilde{b}_{\mu\nu}$ field is invariant under both local Lorentz and gauge transformations. Consequently, the need for anomaly inflow is eliminated: anomalies no longer appear in the effective action, up to covariant field redefinitions. This trivialization suggests that anomaly cancellation is no longer an external consistency condition, in the NR limit of heterotic string theory. In particular, it implies that the anomaly polynomial becomes cohomologically trivial at this order.
III.2 Relaxation of gauge group constraints
Relativistic heterotic string theory admits only specific gauge groups—namely $\mathrm{SO}(32)$ and $E_{8}\times E_{8}$—as required by the factorization of the anomaly polynomial,
$I_{12}\sim\left(\mathrm{tr}F^{2}-\mathrm{tr}R^{2}\right)^{2}.$ (25)
This condition arises from the Green-Schwarz mechanism. In contrast, the NR limit trivializes this mechanism, eliminating the need for such factorization, and potentially removing constraints on the gauge group. While the analysis in this Letter is restricted to the bosonic sector and valid up to four-derivative order in the absence of fermions, it strongly suggests that consistent NR gauge sectors can extend beyond conventional options. This includes $\mathrm{SU}(N)$, products of simple or semi-simple groups, which are excluded in the relativistic setting. However, if the NR theory is understood as a strict limit of a fixed relativistic heterotic model, the gauge group must be fixed demanding compatibility with the relativistic theory (e.g., $\mathrm{SO}(32)$ or $E_{8}\times E_{8}$), and the NR limit inherits this choice. Therefore, the relaxation of gauge group constraints should be viewed as a genuine feature of intrinsically NR theories—that is, theories not obtained as limits of relativistic string backgrounds. From a more general perspective, the structure enabling this relaxation, i.e. the presence of a Chern-Simons-like three-form in the definition of the $H$-flux, appears in a wider class of theories beyond heterotic strings, including gauged supergravities. The trivialization mechanism uncovered here may thus reflect a deeper property of Chern-Simons couplings under NR limits  [^2] We thank to D. Marques for sharing this enlighting observation..
III.3 Modified topological constraints in compactifications
In relativistic string compactifications, anomaly cancellation imposes global topological constraints on the spacetime and gauge bundles. This is encoded in the Bianchi identity for the $H$-flux,
$dH=\mathrm{tr}(F\wedge F)-\mathrm{tr}(R\wedge R),$ (26)
which enforces a matching of second Chern classes GSW2 :
$c_{2}(V)=c_{2}(TX).$ (27)
In the NR setting, the Green-Schwarz transformations are absorbed by a field redefinition of the Kalb-Ramond field: $\tilde{b}=b+D$, where $D$ is the field redefinition which trivializes the gauge and gravitational GS mechanisms, simultaneously. Crucially, the corresponding three-form field strength becomes exact:
$\bar{h}=d\tilde{b}=db+dD=d(b+D),$ (28)
so its Bianchi identity satisfies
$d\bar{h}=0$ (29)
identically, without requiring a nontrivial cancellation between $\mathrm{tr}(F\wedge F)$ and $\mathrm{tr}(R\wedge R)$. As a result, the anomaly inflow condition becomes structurally trivial, and no cohomological constraint on gauge or tangent bundles remains. This opens up the possibility of previously forbidden compactifications including those with non-matching Chern classes, non-Kähler internal geometries, or background torsion in the NR limit.
III.4 Wald entropy and the first law in non-relativistic heterotic backgrounds
The analysis of black hole thermodynamics in the presence of higher-derivative corrections has traditionally relied on Wald’s formalism Wald . For heterotic supergravity at order $\alpha^{\prime}$, the relevant Noether charges require careful treatment due to the Green-Schwarz couplings between the Kalb-Ramond field and the gauge and gravitational Chern-Simons termsOrtin . In the NR limit, the results of this letter predict a dramatic simplification of this structure. Since both the gauge and Lorentz Green-Schwarz mechanisms are trivialized, the redefined Kalb-Ramond field transforms covariantly, and no Chern-Simons modifications are induced. As a result, the Bianchi identity for the corrected three-form curvature reduces to a standard closure condition in the NR limit, as in Eq. (29). This implies that the horizontal Kalb-Ramond momentum map $P_{k}$ becomes simpler, relaxing its role in the entropy computation (with a similar argument we expect a relaxation in the vertical Yang-Mills momentum map). The trivialization of the Green-Schwarz mechanism in the NR limit of heterotic string theory thus enables a simpler derivation of the first law of black hole mechanics, in which all thermodynamic potentials and conserved charges are computed from manifestly gauge- and Lorentz-invariant expressions, up to four-derivative contributions.
III.5 Connection to duality-invariant frameworks
The NR limit of heterotic string theory aligns naturally with duality-invariant formulations such as Double Field Theory (DFT), where generalized geometry encodes both geometric and gauge degrees of freedom EandD . The four-derivative extension of DFT Tduality Hohm involves a generalized vielbein formalism DFTframe . It is therefore not clear whether the NR limit produces a finite four-derivative Lagrangian considering the flux formulation of DFT, compatible with a generalization of the Green-Schwarz mechanism Odd . This motivates a complete NR extension of heterotic DFT, incorporating supersymmetry and $\alpha^{\prime}$ corrections beyond leading order, extending the results of LNR for the relativistic case.
IV Conclusion
We have proved that the Green-Schwarz mechanism of heterotic string theory can be trivialized in the NR limit up to four-derivative terms. We constructed the field redefinition of the b-field that makes this possible and, by construction, the Bianchi identity for the 3-form curvature does not include $\alpha^{\prime}$-corrections. Our results establish a new paradigm for anomaly cancellation in the NR limit of heterotic string theory with several implications and open directions for future research in the long term. The immediate continuation of this work includes:
  • Investigating the role of supersymmetry when higher-derivative terms are considered, extending the seminal work in Bsusy and exploring further trivializations of the supersymmetric transformations at the NR supergravity level Bdroo .
  • Understanding the embedding of the duality invariant construction of LNR for the higher-derivative structure of heterotic DFT in the NR limit, including the construction of the full four-derivative heterotic finite action (including fermions) in the NR limit.
  • Constructing the simplified first law and Wald entropy formula of NR heterotic stringy black holes at first order in $\alpha^{\prime}$.
The previous ideas follow a clear path towards the understanding of the $\alpha^{\prime}$- corrected heterotic supergravity under the NR limit.
Acknowledgements.
E.L. is very grateful to Diego Marques for discussions. This work is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.